• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 22:45:25 on August 21, 2009

    Entry number 285179

    Followups:

    keyword=Locking cavity after EPICS reboot painful

    The cavity would not lock after the EPICS reboot.

    After another EPICS reboot I lost totally control of the laser and
    rebooting EPICS would not help.

    The empirical fix to the problem is to reboot the Compton computer and
    then do an EPICS reboot.

    I got control back of the laser after rebooting EPICS after the computer
    reboot.

    One detail to know is the EPICS variable
    COMPTON_LASERT_MIN has to be set to 35.8 instead of the default value 32.2

    Not sure why we had to do that, the only thing I can see possible is the
    Compton computer polling to much on the IOC for unknown reasons ( maybe
    network sockets left opened ). To be investigated...




    A copy of this log entry has been emailed to: nanda@jlab.org,withers@jlab.org,mdalton@jlab.org